EPS-HEP99 Abstract # 000 Parallel sessions: Plenary sessions:

Study of Three-Jet Production in Deep-Inelastic Positron Proton Collisions at HERA

H1 Collaboration,

DESY, Notkestr.85, D-22607 Hamburg, Germany

The production of three-jet events has been studied in deep-inelastic positron proton collisions using data collected with the H1 detector in 1995-1997. In the kinematic range of four-momentum transfers 5 < Q^{2} < 15000 GeV^{2} a set of variables describing the three-jet system is investigated. The cross sections as a function of the invariant three-jet mass M_{3J}, the average transverse jet energy in the Breit frame \overline{E}_{T,Breit} as well as x_{Bj} and Q^{2} are measured. Furthermore the shape of the Dalitz variables X_{3} = E_{3}/M_{3J}, X_{4} = E_{4}/M_{3J}, where E_{3} (E_{4}) is the energy of the (second) highest jet in the three jet center of mass frame, and two angular distributions are studied. Events with 3-jet masses up to M_{3J} = 130 GeV and \overline{E}_{T,Breit} = 30 GeV have been found in the data sample. The results are compared to the predictions of parton cascade models and to perturbative calculations in O(α_{s}^{2}).